OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -one halfx + 5. y = -2x + 2 y = 2x - 14 y = one halfx - 8 y = -one halfx - 4

OpenStudy (anonymous):

in order for lines to be parallel they must have the same slope

OpenStudy (anonymous):

the standard equation of a line is y=mx+b

OpenStudy (anonymous):

the line you are given it y=-(1/2)x+5

OpenStudy (anonymous):

so you can see m=-1/2

OpenStudy (anonymous):

so you have y=-(1/2)x+b so far for your answer

OpenStudy (anonymous):

you are told it must go through (4,-6) so you know when x=4 y=-6

OpenStudy (anonymous):

so you can use that to solve for b

OpenStudy (anonymous):

-6=(-1/2)(4)+b

OpenStudy (anonymous):

So its D?

OpenStudy (anonymous):

y = -1/2x - 4

OpenStudy (anonymous):

yup